OpenStudy (anonymous):

Someone helppp? The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

so what is the line equation for y-4=2/3(x-3)?

OpenStudy (anonymous):

are you able to solve for y?

OpenStudy (anonymous):

What formula do you use?

OpenStudy (anonymous):

there's no formula

OpenStudy (anonymous):

Then how do i solve?

OpenStudy (anonymous):

k let start over. what's your slope?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

so parallel line you will have a same slope but different b right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then use y-y1=2/3(x-x1)

OpenStudy (anonymous):